OpenStudy (hy123):

(2(sqrt)7-3)/(sqrt)(7)-2) rationalize the denominator and simplify

OpenStudy (hy123):

\[2\sqrt{7}-3/\sqrt{7}-2\]

OpenStudy (hy123):

Hi Mukushla, I got \[8+\sqrt{7}/5\]

OpenStudy (hy123):

Is that correct?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

completely right u did it well

OpenStudy (hy123):

Thanks :D

OpenStudy (hy123):

mukusha since you are already here mind helping me with another rpoblem?

OpenStudy (hy123):

s(x)=x+1 t(x)=2x^2

OpenStudy (hy123):

(s of t)(x)=2x^2+1?

OpenStudy (anonymous):

yes

OpenStudy (hy123):

Well it says (s x t)(x)

OpenStudy (anonymous):

(s x t)(x)=s(x)*t(x)

OpenStudy (hy123):

So (x+1)(2x^2)?

OpenStudy (anonymous):

thats right

OpenStudy (hy123):

ty